OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −1/3x − 6 y = −1/3x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (mysesshou):

y=3x+3 m=3 parallel m = 3 point (4, −2). (y-y1) = m* (x-x1) y - (-2) = 3 * (x-4) simplify.

OpenStudy (apoorvk):

B. this question was posted earlier.. check out http://openstudy.com/study#/updates/4f6a5a61e4b0ba795ab39da4